OpenStudy (anonymous):

12. Solve x/3 - 1 = -4 for x. (Points : 1) -9 -3 3 9 13. Solve 22x - 4 = 40 for x. (Points : 1) 1 2 4 10 14. Solve x/4 + 3 = 1 for x. (Points : 1) 12 8 -12 -8

OpenStudy (lyrae):

x/3 - 1 = -4 <=> x/3 = -3 <=> x = -9 22x - 4 = 40 <=> 22x = 44 <=> x = 2 x/4 + 3 = 1 <=> x/4 = -2 <=> x = -8